OpenStudy (anonymous):

Christy drove 300 miles on her vacation. She drove an average of 1.25 times faster on the second 150 miles of her trip. Which expression represents the time she spent driving? Let x= her speed on the first half of the trip. a.375/x b. 337.5/x c.270x d. 270/x

OpenStudy (anonymous):

I had the same question and the answer was C.

OpenStudy (anonymous):

\[d=r \times t \] so \[\frac{ d }{ r}= t\] Does that help?

OpenStudy (anonymous):

\[150/x=\] time for the first have and since 1.25x is the rate for the second half of the trip then \[\frac{ 150 }{ 1.25x } = t\] for the second half of the trip

OpenStudy (anonymous):

then simply complete the operations...

OpenStudy (anonymous):

i.e. add the two times to get a total time...

OpenStudy (anonymous):

hope that helped